A police car drives past you with its siren on. While the car is approaching, you hear a frequency of 550 Hz. A Recalling that the speed of sound is 767 mph, what is the speed of the police car if the siren has a frequency of 507 Hz when neither you nor the car are moving? B What would the apparent frequency of the siren be if you were driving toward the parked police car at this same speed.

A.60mph

b.547Hz
Answer #1

A)while source of sound is approaching observer

Apparent frequency=550 Hz

Speed of sound=767mph

Let speed of car be v

Actual frequency of sound=507 mph

550=507*(767/767-v)

v=60 mph

B)

Apparent frequency=(767+60)/767 * 507=547 Hz